# A Branch-and-Cut-and-Price Algorithm for Cutting Stock and Related Problems

### v2

@@ 9.3 Ordered Open-End Bin Packing Problem
Let $\mathcal{P}_{i}$ be the set of patterns where $i$ is the last item packed, and $\gamma_{i}$ denote the usage of item $i$ as the last item, i.e., $\gamma_{i}=\sum_{P\in P_{i}}\lambda_{P}$. If there exists any item $i$ such that $10<\gamma_{i}<$ 1, we proceed to branch on the item with a value closer to $1$. The left branch considers only patterns where $i$ is the last item, while the right branch considers only patterns where $i$ is not the last item. The pricer efficiently manages this branching scheme by excluding item $i$ from dynamic programming on the left branch and not considering item $i$ as the last item during pattern recovery on the right branch. If $\gamma_{i}=0$ or $\gamma_{i}\geq 1$ (where $\gamma_{i}>1$ indicates overpacking of item $i$) for all items, we resort to branching using the aforementioned Ryan-Foster scheme.
@@ 9.4 Class-Constrained Bin Packing Problem
where items in $I$ are sorted in non-decreasing order of $q_{i}$ and $i$ is the index of the current item. The variables $r$, $c$, and $u$ represent the remaining capacity, remaining permissible class types, and a binary indicator, respectively. Additionally, $\overline{u}$ signifies the complement of $u$, while $e$ indicates whether the classes of items $I[i]$ and $I[i-1]$ are equal. The lower bound for $c_{P}(\overline{\pi},\overline{\rho},1)$ corresponds to the value of the state $f(n,W,C,0)$.
@@ Acknowledgments
Supported by the São Paulo Research Foundation (FAPESP) grants #2015/11937-9 and #2022/05803-3; and the Brazilian National Council For Scientific and Technological Development (CNPq) grant #312345/2023-2.
@@ Appendix A Proof of the Conflict Propagation’s Lemma
Let $v(i,j)$ denote the current B&B node, where the Extended Ryan-Foster Scheme selects the pair $(i,j)$ of superitems for branching. Let $k$ represent the item such that $w_{k}=w_{i}+w_{j}$. Additionally, node $v(i,j)$ possesses a set of superitems $\overline{I}$, where each superitem $i$ is a nonempty subset $S\subseteq I$ (the items from the original instance). Alongside, there exists a conflict graph $G=(\overline{I},E)$, where each edge $\{k,l\}\in E$ represents a conflict introduced by a right branch of our branching scheme. We denote by $E_{i}$ the set of superitems that have a conflict with superitem $i$ in $G$, i.e., the adjacency list of vertex $i$ in $G$.
An alternative approach to Definition 2 is to consider the edges added to $G$ rather than focusing solely on modifications in the conflict lists. Let $G=(\overline{I},E)$ and $G^{\prime}=(\overline{I}^{\prime},E^{\prime})$ represent the conflict graphs before and after branching at the current node $v(i,j)$. In this context, the left branch defines $E^{\prime}=E\cup\{\{a,k\}\colon\{a,i\}\in E\vee\{a,j\}\in E\}$, while the right branch defines $E^{\prime}=E\cup\{\{i,j\}\}$.
@@
Let $v(k^{\prime},h)$ be the ancestor of $v(i,j)$ that added the conflict $\{k^{\prime},h\}$ in its right branch. Let $P_{2}$ be a pattern containing item $k^{\prime}$ in $\overline{\lambda}$, and let $\overline{\lambda}^{\prime}$ be the equivalent solution obtained by swapping $k^{\prime}$ and $k$, with $P_{1}^{\prime}=P_{1}\cup\{k^{\prime}\}\setminus\{k\}$ and $P_{2}^{\prime}=P_{2}\cup\{k\}\setminus\{k^{\prime}\}$. If there is no conflict in $P_{1}^{\prime}$ and $P_{2}^{\prime}$, then $\overline{\lambda}^{\prime}$ is an improvement solution in the search space of the left child of $v(k^{\prime},h)$, which leads to a contradiction (as we prioritize left in the DFS order).
Suppose, on the contrary, that either there is a conflict $\{k^{\prime},g\}$ in $P_{1}^{\prime}$, with $g\in E_{k^{\prime}}$, or there is a conflict $\{i,g\}$ or $\{j,g\}$ in $P_{2}^{\prime}$, with $g\in E_{i}\cup E_{j}$. Without loss of generality, as all cases are analogous, let $\{k^{\prime},g\}$ be the earliest conflict. Observe that this conflict was introduced by the ancestor $v(k^{\prime},g)$ of $v(k^{\prime},h)$ in its right branch. Thus, $\overline{\lambda}^{\prime}$ is a feasible and improvement solution in the left child of $v(k^{\prime},g)$, which is, again, a contradiction.   $\blacksquare$
@@
Proof: Clearly, $E_{k^{\prime}}\subseteq E_{k^{\prime}}^{\prime}$, as by the branching scheme, conflicts are inherited. Now, it is invalid to define $E_{i}\cup E_{j}\subseteq E_{k^{\prime}}^{\prime}$ only if there exists an item $h\in E_{i}\cup E_{j}$ such that there is an improvement solution $\overline{\lambda}$ in the subtree of $v(i,j)$ containing a pattern $P_{1}$ with $h$ and $k^{\prime}$ together.

### v3

@@ 9.3 Ordered Open-End Bin Packing Problem
Let $\mathcal{P}_{i}$ be the set of patterns where $i$ is the last item packed, and $\sigma_{i}$ denote the usage of item $i$ as the last item, i.e., $\sigma_{i}=\sum_{P\in\mathcal{P}_{i}}\lambda_{P}$. If there exists any item $i$ such that $10<\sigma_{i}<$ 1, we proceed to branch on the item with a value closer to $1$. The left branch considers only patterns where $i$ is the last item, while the right branch considers only patterns where $i$ is not the last item. The pricer efficiently manages this branching scheme by excluding item $i$ from dynamic programming on the left branch and not considering item $i$ as the last item during pattern recovery on the right branch. If $\sigma_{i}=0$ or $\sigma_{i}\geq 1$ (where $\sigma_{i}>1$ indicates overpacking of item $i$) for all items, we resort to branching using the aforementioned Ryan-Foster scheme.
@@ 9.4 Class-Constrained Bin Packing Problem
where items in $I$ are sorted in non-decreasing order of $q_{i}$ and $i$ is the index of the current item. The variables $r$ and $c$ represent, respectively, the remaining capacity and number of classes. Additionally, $u$ and its complement $\overline{u}$ are binary values, where $u$ indicates if $q_{I[i]}$ was already discounted from $c$, while $e$ indicates whether $q_{I[i]}$ and $q_{I[i-1]}$ are equal. The lower bound for $c_{P}(\overline{\pi},\overline{\rho},1)$ corresponds to the value of $f(n,W,C,0)$.
@@ Acknowledgments
Supported by the São Paulo Research Foundation (FAPESP) grants #2015/11937-9 and #2022/05803-3; the Brazilian National Council For Scientific and Technological Development (CNPq) grant #312345/2023-2; and Teaching, Research and Extension Support Fund of University of Campinas (FAEPEX/UNICAMP) grant #2372/23.
@@ Appendix A Proof of the Conflict Propagation’s Lemma
Let $v(i,j)$ denote the current B&B node, where the Extended Ryan-Foster Scheme selects the pair $(i,j)$ of superitems for branching. Let $k$ represent the item such that $w_{k}=w_{i}+w_{j}$. Additionally, node $v(i,j)$ possesses a set of superitems $\overline{I}$, where each superitem $i$ is a nonempty subset $S\subseteq I$ (the items from the original instance). Alongside, there exists a conflict graph $G=(\overline{I},E)$, where each edge $\{k,l\}\in E$ represents a conflict introduced by a right branch of our branching scheme or by conflict propagation. We denote by $E_{i}$ the set of superitems that have a conflict with superitem $i$ in $G$, i.e., the adjacency list of vertex $i$ in $G$.
An alternative approach to Definition 2 is to consider the edges added to $G$ rather than focusing solely on modifications in the conflict lists. Let $G=(\overline{I},E)$ and $G^{\prime}=(\overline{I}^{\prime},E^{\prime})$ represent the conflict graphs before and after branching at the current node $v(i,j)$. In this context, the left branch defines $E^{\prime}=E\cup\{\{a,k\}\colon\{a,i\}\in E\vee\{a,j\}\in E\}$, while the right branch defines $E^{\prime}=E\cup\{\{i,j\}\}$. An interesting observation is that the conflict added to a node $v(i,j)$ remains valid for all its descendants, regardless of whether the item demand becomes zero at some point.
@@
Let $v(k^{\prime},h)$ be the ancestor of $v(i,j)$ that added the conflict $\{k^{\prime},h\}$ in its right branch. Let $P_{2}$ be a pattern containing item $k^{\prime}$ in $\overline{\lambda}$, and let $\overline{\lambda}^{\prime}$ be the equivalent solution obtained by swapping $k^{\prime}$ and $k$, with $P_{1}^{\prime}=P_{1}\cup\{k^{\prime}\}\setminus\{k\}$ and $P_{2}^{\prime}=P_{2}\cup\{k\}\setminus\{k^{\prime}\}$. It is possible that $k^{\prime}$ does not exist as an individual item because it was merged with other items to form a superitem $l$. In this case, we redefine $l$ to be created using $k$ instead of $k^{\prime}$. If there is no conflict in $P_{1}^{\prime}$ and $P_{2}^{\prime}$, then $\overline{\lambda}^{\prime}$ is an improvement solution in the search space of the left child of $v(k^{\prime},h)$, which leads to a contradiction (as we prioritize left in the DFS order).
Suppose, on the contrary, that either there is a conflict $\{k^{\prime},g\}$ in $P_{1}^{\prime}$, with $g\in E_{k^{\prime}}$, or there is a conflict $\{i,g\}$ or $\{j,g\}$ in $P_{2}^{\prime}$, with $g\in E_{i}\cup E_{j}$. The conflicts $\{i,g\}$ and $\{j,g\}$ might occur inclusive within the superitem $l$, but it does not affect our argument. Without loss of generality, as all cases are analogous, let $\{k^{\prime},g\}$ be the earliest conflict. Observe that this conflict was introduced by the ancestor $v(k^{\prime},g)$ of $v(k^{\prime},h)$ in its right branch. Thus, $\overline{\lambda}^{\prime}$ is a feasible and improvement solution in the left child of $v(k^{\prime},g)$, which is, again, a contradiction.   $\blacksquare$
@@
Proof: Clearly, $E_{k^{\prime}}\subseteq E_{k^{\prime}}^{\prime}$, as by the branching scheme, conflicts are inherited. Now, it is invalid to define $E_{i}\cup E_{j}\subseteq E_{k^{\prime}}^{\prime}$ only if there exists an item $h\in E_{i}\cup E_{j}$ such that there is an improvement solution $\overline{\lambda}$ in the subtree of $v(i,j)$ containing a pattern $P_{1}$ with $h$ and $k^{\prime}$ together. As said before, if $k^{\prime}$ was used to create a superitem $l$ at $v(i,j)$, redefine $l$ to be created with $k$.